OpenStudy (anonymous):

What is the equation of a hyperbola with a=6 and c=8? Assume that the transverse axis is horizontal.

OpenStudy (alexwee123):

do we also assume that the hyperbola is centered at the origin?

OpenStudy (anonymous):

im assuming

OpenStudy (valpey):

Is 'a' something like the difference between the foci and 'c' the common difference?

OpenStudy (anonymous):

it doesnt say.

OpenStudy (alexwee123):

k then you would find the value of "b" with c^2=a^2 + b^2 64=36+b^2 b=squareroot 28

OpenStudy (alexwee123):

and 'a' means semi length of the tranverse axis for this problem @Valpey :O

OpenStudy (alexwee123):

then because the transverse axis is horizontal then we would plug in the values of "a" and "b" into the hyperbola equation